so the equaiton is: measure of angle 2=x+78. The triangles point is 55 degree and thr right side and bottom is the same length. How do I solve this?
Answer by stanbon(75887)

so the equaiton is: measure of angle 2=x+78. The triangles point is 55 degree and thr right side and bottom is the same length. How do I solve this?
------
If 2 sides are the same length, the angles opposite those sides are equal.
-----
You are given that the vertex angle is 55 degrees.
----
The sum of the 2 base angles is 180-55 = 125 degrees
----
Each base angle is 125/2 = 62.5 degrees
------
Note: I don't know where your #2 angle is.
-------------------------------------------
Cheers,
Stan H.